OpenStudy (anonymous):

Solve the equation. If it is an identity (all real numbers) or a contradiction (no solution), so state. 4x + 3 + 3x = 8x + 3 – x

OpenStudy (anonymous):

it is an identity (all real numbers)

OpenStudy (anonymous):

real real values of x will satisfy that. SO its solution is all real numbers

OpenStudy (anonymous):

all real values...

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

4x+3+3x=8X+3-x 7x+3=7x+3 7x-7x=3-3=0